OpenStudy (anonymous):

Find 3 consecutive odd numbers where the product of the smaller two numbers is 34 less than the square of the largest number.

OpenStudy (anonymous):

@blackbird02

OpenStudy (anonymous):

@sid0830

OpenStudy (anonymous):

Let the three odd numbers be: x, x+2 and x+4 According the the question, the equation is formed as below: (x)(x+2)=(x+4)^2-34 Solving this, we get x=3 Other two numbers are 5 and 7

OpenStudy (anonymous):

See, if we take difference between two odd number is minimum 2. Because if we will take difference of 1, the next number will be an even number but we want all the tree odd numbers. That's why I used x+2..!!